OpenStudy (anonymous):

what is the total money collected, t, varies directly with the number of student tickets sold. Write the equation of direct variation if the school collected $1107 for 246 student tickets at the basketball game?

OpenStudy (anonymous):

since t is directly proportional to the number of students t = k x no. of students k = 1107/246 = 9/2 t = 9/2 x no. of students

OpenStudy (anonymous):

so t= 20.25?

OpenStudy (anonymous):

noo... t and no. of students is given... what matters is k because it shows the ratio of t to no. of students

OpenStudy (anonymous):

well you just said divide the 1107 by the number of students and that gives you 4.5, then you said mult. 9/2 by the number of students?

OpenStudy (anonymous):

so the equation is t=(9/2)x 246?

OpenStudy (anonymous):

mmm think of it this way in line equation y = mx +c we find out the m and c and substitue the numbers, yet we leave y and x as it is for coordinates.

OpenStudy (anonymous):

its t = 4.5 x no. of students

OpenStudy (anonymous):

oh so i had the 4.5 part right?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

hahah okay thanks!

OpenStudy (anonymous):

sure ^^